# Optimal Look-back Horizon for Time Series Forecasting in Federated Learning

### v1

@@ Introduction
$\begin{split}X_{t,k}^{(H)}&=\big[x^{(k)}_{t-H+1},\dots,x^{(k)}_{t}\big]\in\mathbb{R}^{F\times H},\\ Y_{t,k}^{(S)}&=\big[x^{(k)}_{t+1},\dots,x^{(k)}_{t+S}\big]\in\mathbb{R}^{F\times S}.\end{split}$ (1)
@@
$\begin{split}g_{k}(H)&=2\sum_{j=1}^{J}w_{j,k}\cdot\min\left(1,\frac{H}{T_{j,k}^{*}}\right),\\ w_{j,k}&=\frac{\sum_{f=1}^{F}A_{f,j,k}^{2}}{\sum_{f=1}^{F}\sum_{j=1}^{J}A_{f,j,k}^{2}}.\end{split}$ (6)
@@
$\begin{split}L_{\mathrm{AR}}^{(k)}(S)&:=\mathbb{E}\big[\|\ Y_{\mathrm{AR},k}^{(S)}-\mathbb{E}[Y_{\mathrm{AR},k}^{(S)}\mid X_{k}^{(H)}]\ \|_{2}^{2}\big],\\ L_{\mathrm{seas}}^{(k)}(H)&:=\mathbb{E}\big[\|\ Y_{\mathrm{seas},k}^{(S)}-\mathbb{E}[Y_{\mathrm{seas},k}^{(S)}\mid X_{k}^{(H)}]\ \|_{2}^{2}\big],\\ L_{\mathrm{trend}}^{(k)}(H)&:=\mathbb{E}\big[\|\ Y_{\mathrm{trend},k}^{(S)}-\mathbb{E}[Y_{\mathrm{trend},k}^{(S)}\mid X_{k}^{(H)}]\ \|_{2}^{2}\big].\end{split}$ (13)
@@ Acknowledgments
    $\mu_{f,k}=\frac{1}{H}\sum_{i=1}^{H}x_{f,i,k},\quad\sigma_{f,k}=\sqrt{\frac{1}{H}\sum_{i=1}^{H}(x_{f,i,k}-\mu_{f,k})^{2}}$ (36)
@@
    $\begin{split}g_{k}(H)&=2\sum_{j=1}^{J}w_{j,k}\,\min\!\Big(1,\frac{H}{T_{j,k}^{*}}\Big)\,,\\ w_{j,k}&=\frac{\sum_{f=1}^{F}A_{f,j,k}^{2}}{\sum_{f=1}^{F}\sum_{j=1}^{J}A_{f,j,k}^{2}}\end{split}$ (42)
@@
$\begin{split}L_{\mathrm{Bayes}}&:=\mathbb{E}[\|Y-m^{*}(X)\|^{2}],\\ L_{\mathrm{approx}}&:=\mathbb{E}[\|m^{*}(X)-m(X)\|^{2}].\end{split}$ (58)

### v2

@@ Abstract
Extended Version — The full version is provided on arXiv.
@@ Introduction
$X_{t,k}^{(H)}=\big[x^{(k)}_{t-H+1},\dots,x^{(k)}_{t}\big]\in\mathbb{R}^{F\times H},$ (1)
$Y_{t,k}^{(S)}=\big[x^{(k)}_{t+1},\dots,x^{(k)}_{t+S}\big]\in\mathbb{R}^{F\times S}.$ (2)
@@
$g_{k}(H)=2\sum_{j=1}^{J}w_{j,k}\cdot\min\left(1,\frac{H}{T_{j,k}^{*}}\right),$ (7)
$w_{j,k}=\frac{\sum_{f=1}^{F}A_{f,j,k}^{2}}{\sum_{f=1}^{F}\sum_{j=1}^{J}A_{f,j,k}^{2}}.$ (8)
@@
$L_{\mathrm{AR}}^{(k)}(S):=\mathbb{E}\big[\|\ Y_{\mathrm{AR},k}^{(S)}-\mathbb{E}[Y_{\mathrm{AR},k}^{(S)}\mid X_{k}^{(H)}]\ \|_{2}^{2}\big],$ (15)
$L_{\mathrm{seas}}^{(k)}(H):=\mathbb{E}\big[\|\ Y_{\mathrm{seas},k}^{(S)}-\mathbb{E}[Y_{\mathrm{seas},k}^{(S)}\mid X_{k}^{(H)}]\ \|_{2}^{2}\big],$ (16)
$L_{\mathrm{trend}}^{(k)}(H):=\mathbb{E}\big[\|\ Y_{\mathrm{trend},k}^{(S)}-\mathbb{E}[Y_{\mathrm{trend},k}^{(S)}\mid X_{k}^{(H)}]\ \|_{2}^{2}\big].$ (17)
@@ Acknowledgments
    $\mu_{f,k}=\frac{1}{H}\sum_{i=1}^{H}x_{f,i,k},$ (40)
    $\sigma_{f,k}=\sqrt{\frac{1}{H}\sum_{i=1}^{H}(x_{f,i,k}-\mu_{f,k})^{2}}$ (41)
@@
    $g_{k}(H)=2\sum_{j=1}^{J}w_{j,k}\,\min\!\Big(1,\frac{H}{T_{j,k}^{*}}\Big)$ (47)
    $w_{j,k}=\frac{\sum_{f=1}^{F}A_{f,j,k}^{2}}{\sum_{f=1}^{F}\sum_{j=1}^{J}A_{f,j,k}^{2}}$ (48)
@@
$L_{\mathrm{Bayes}}:=\mathbb{E}[\|Y-m^{*}(X)\|^{2}],$ (64)
$L_{\mathrm{approx}}:=\mathbb{E}[\|m^{*}(X)-m(X)\|^{2}].$ (65)